Please help! I don't understand! The problem is:
y=6x-4
y=-2x+28

Answer by ankor@dixie-net.com(22740)

y=6x-4
y=-2x+28
:
Well, you know that y = y, therefore
6x - 4 = -2x + 28
6x + 2x = 28 + 4
8x = 32
x = 32/8
x = 4
:
Find y
y = 6(4) - 4
y = 24 - 4
y = 20
:
check solutions in the original 2nd equation
y=-2x+28
20 = -2(4) + 28
20 = -8 + 28; confirms our solutions
;
Do you understand it now?